OpenStudy (chrisneedshelp):

Solve x2 - 8x + 3 = 0 by completing the square. Which equation is used in the process? (x - 4)2 = 0 (x - 4)2 = 13 (x - 4)2 = 19

OpenStudy (chrisneedshelp):

@d3Xter

OpenStudy (wolf1728):

Here's a page that will explain completing the square: http://1728.org/quadr2.htm (The second half of the page explains about deriving the quadratic formula so just ignore that). Completing the square: x2 - 8x + 3 = 0 Move the "non X" term to the right x² -8x = -3 Divide the equation by the coefficient of X² That doesn't have to be done because there is no coefficient. get the coefficient of x divide by 2 square it add it to both sides of the equation -8/2 = -4 -4² = 16 x² -8x + 16 = -3 +16 Take the square root of both sides: (x-4) = sqrt(13) x = sqrt (13) + 4 x1 = 3.6055512755 + 4 x1 = 7.6055512755 x2 = -sqrt(13) + 4 x2 = -3.6055512755 + 4 x2 = 0.3944487245

OpenStudy (chrisneedshelp):

thank you !!

OpenStudy (wolf1728):

u r welcome glad to help.